OpenStudy (anonymous):

Let f (x) = 2x2 – 7x – 20 and g(x) = 3x – 1. (a) Find the difference (f – g)(x) and simplify the results. Show work. (b) Find the composite function (fog)(x) and simplify the results. Show work. (c) Find (fog)(-1). Show some work.

OpenStudy (ajprincess):

(f-g)(x)=f(x)-g(x) =2x^2-7x-20-(3x-1) Can u find (f-g)(x) @Rose1285?